OpenStudy (anonymous):

Three times the sum of a number and seven is two less than five times the number. a. 3x+7=5x-2 b. 3(x+7)=5x-2 c. 3(x+7)=2-5x d. 3x+7=2-5x

OpenStudy (anonymous):

CAN U HELP ME EAT MY PUSSY PLEASE!!!

OpenStudy (anonymous):

Negative^ No thanks. Can someone please help me with this problem?

OpenStudy (swissgirl):

Ok so its 3*(the sum of a number and 7) is means equal 2 less from 5*the number So let the number =x So 3*(x+7)= 2 less from 5x

OpenStudy (swissgirl):

Now 2 less from 5x is the tricky part can you figure it out?

OpenStudy (anonymous):

23?

OpenStudy (swissgirl):

23?

OpenStudy (texaschic101):

its B

OpenStudy (swissgirl):

He is right

OpenStudy (texaschic101):

do you need it worked out ?

OpenStudy (swissgirl):

He needs to understand how to translate words into an equation into mathematical terms

OpenStudy (texaschic101):

ok....3 times the sum of a number and 7......the um of a number and 7 = (x + 7) and 3 times this is 3(x + 7)...."is " means equals. Two less then 5 times the number.....5x - 2. Which comes down to 3(x + 7) = 5x - 2 Do you understand ?

OpenStudy (anonymous):

ohhhh yes!!! Thank you guys soooo much!!!

OpenStudy (texaschic101):

YW :)